# Asymptotically Optimal Lazy Lifelong Sampling-based Algorithm for Efficient Motion Planning in Dynamic Environments

### v2

@@ 5       $V\xleftarrow{+}\{v_{new}\}$;
8       if local trajectory exists between $v$ and $u$ then
9             $E\xleftarrow{+}\{e(v_{new},u)\}$;
@@ 10             UpdateNode($v,u$);
12      if local trajectory exists between $u$ and $v$ then
13             $E\xleftarrow{+}\{e(u,v_{new})\}$;
14
15
16
@@ III-B Informed Rewiring Cascade
LLPT dynamically adjusts the solution path by rewiring $\mathcal{T}$ when the edge weights of $\mathcal{G}$ change. Drawing inspiration from RRTX [15] and RRT [7], LLPT employs two cost-to-goal estimates, g($v$) and lmc($v$), for each node to detect local inconsistencies arising from graph modifications. A node $v$ is called consistent if g($v$) = lmc($v$)<absent<$\infty$< ∞, and inconsistent otherwise. An inconsistent node is further categorized as locally overconsistent if lmc($v$) $<$ g($v$), or locally underconsistent if lmc($v$) $>$ g($v$) or lmc($v$) = g($v$) $\infty$. The g-value is the accumulated cost-to-goal by traversing the previous search tree, while the lmc-value is the cost-to-goal based on the up-to-date graph topology and is potentially better informed than the g-value. The lmc-value satisfies the following relationship when all nodes are locally consistent:
@@
Modifications on edge weights potentially trigger changes in the lmc-values of some nodes and render these nodes inconsistent. LLPT collects and stores the inconsistent nodes in a priority queue $\mathcal{Q}$. The priority of each node in the queue is based on the order of their associated keys. If a node $v$ becomes underconsistent, which happens when it solution path to goal is evaluated to be in collision, LLPT will use the PropagateCostToLeave procedure (Algorithm 3) to propagate the cost ascendance to its leaf nodes. Afterwards, both $v$ and its leaf nodes will be added to the priority queue $\mathcal{Q}$. On the other hand, if a node $v$ becomes overconsistent, which can happen when some of its outgoing edges have decreasing weight or when $\mathcal{G}$ is extended, $v$ will be added to $\mathcal{Q}$.
The ComputeShortestPath procedure, which is provided in Algorithm 4, is called to make the spanning tree consistent again by operating on the inconsistent nodes stored in $\mathcal{Q}$ iteratively. For a underconsistent node, ComputeShortestPath iteratively searches among its neighbor nodes, and the underconsistent node is rewired to a new parent node if the new parent node resulting in a better cost-to-goal (line 4-4, Algorithm 4). For a overconsistent node, its local inconsistency is disseminated to its neighbors. Specifically, if $v$ emerges as a better parent node for a neighbor than its current parent, the neighbor’s parent is updated to $v$, and the neighbor will be included in $\mathcal{Q}$ for further processing until consistency is restored within the nodes relevant to the solution (line 4-4, Algorithm 4).
@@
where $\dot{<}$ is a lexicographical comparator with $k(v)\dot{<}k(u)$ if either $k_{1}(v)<k_{2}(u)$ or $k_{1}(v)=k_{1}(u)\land k_{1}(\mathbf{v})<k_{2}(u)$. The node in $\mathcal{Q}$ with the smallest key is regarded as the most promising one and will be popped out from $\mathcal{Q}$ and processed first until there is not node with smaller key than $v_{start}$ and $v_{start}$ has found a solution path to $v_{goal}$ (line 1, Algorithm 4). Since $v_{start}$ is updated as the robot moves, $h(v_{start},v)$ is changing and thus the key used by the priority queue. To avoid having to reorder the queue, we adopt the key updating strategy as D Lite[5], where a constant $k_{m}$ (initialized as 0) that is added up with the cost of the robot movement at each replanning iteration is used to mitigate the priority inconsistency of the out-of-date priority in $\mathcal{Q}$ with a newly-inserted vertex (line 1, Algorithm 1). The key of a newly-inserted vertex $k(v)$ is calculated by $k_{1}(v)=\min(lmc(v),g(v))+h(v_{start},v)+k_{m}$ and $k_{2}(v)=\min(lmc(v),g(v))$. For each node popped from the top of $\mathcal{Q}$, its key will be recalculated, and the node will be reinserted into $\mathcal{Q}$ with updated priority if its old key is smaller than the new one (line 4-4, Algorithm 4).
@@ III-C Algorithmic framework
The proposed algorithm is provided in Algorithm 1 with its major sub-routines outlined in Algorithms 2-6. Notations such as $A\xleftarrow{+}B$ represents $A\leftarrow A\cup B$, and $A\xleftarrow{-}B$ represents $A\leftarrow A\backslash B$. Set cardinality is denoted by ||$\cdot$|| ⋅ |. $\textnormal{{Child}}(v)$ and $\textnormal{{Parent}}(v)$ retrieve child nodes and the parent of node $v$ in $\mathcal{T}$, respectively. $N(v)$ returns the neighboring nodes of $v$ in $\mathcal{G}$. The priority queue $\mathcal{Q}$ stores nodes sorted in ascending order of their keys. $\mathcal{Q}.\textnormal{{update}}(v)$ insert or re-order node $v$ based on its key value. $\mathcal{Q}.\textnormal{{pop}}()$ pops and returns the vertex with the highest priority in $\mathcal{Q}$. $\mathcal{Q}.\textnormal{{topkey}}$ returns the key of the vertex with the highest priority in $\mathcal{Q}$. The function $\textnormal{{MakeParentOf}}(v,u)$ establishes a child-parent relationship between $v,u$ and adds $v$ to $\mathcal{T}$. $\textnormal{{RemoveFromTree}}(v)$ eliminates the child-parent relationship between $v$ and its parent node and removes $v$ from $\mathcal{T}$. The comparison operator with $\infty$ is defined such that $x<\infty$ is true for any $x\neq\infty$, otherwise false.
@@
LLPT initiates by setting the goal node $v_{goal}$ with lmc($v_{goal}$)=g($v_{goal}$)=0 and the start node $v_{start}$ with lmc($v_{start}$)=g($v_{start}$)$\infty$. The replanning procedure (line 1-1, Algorithm 1) iterates indefinitely at a user-defined frequency until the robot reaches the goal configuration. At the start of each replanning cycle, the search graph $\mathcal{G}$ is refreshed with lazily-computed edges (line 1, Algorithm 1). For edges with updated weights, their starting nodes are rewired using UpdateNode (line 1-1, Algorithm 1). Each replanning cycle comprises two primary loops: the outer loop (line 1-1, Algorithm 1) and the inner loop (line 1-1, Algorithm 1). The inner loop serves as the main search process, rewiring the search tree $\mathcal{T}$ and identifying the shortest collision-free path. This loop alternates between rewiring $\mathcal{T}$ to compute the shortest path and evaluating the edges along the shortest path until a collision-free shortest path is discovered. The procedure EvaluateEdge evaluates the first $\alpha$ unevaluated edges of the shortest path, where $\alpha\in\mathbb{N}^{*}$ is a user-specified parameter to balance the competing computational costs of edge evaluation and rewiring, and returns the collided nodes. For the collided nodes, their lmc-value is increased to infinity. Then, they are removed from $\mathcal{T}$ and inserted into $\mathcal{Q}$ (line 1-1, Algorithm 1). The cost-to-goal increasing information is propagated to their descendants via the procedure PropagateCostToLeave (line 1, Algorithm 1). The inner loop then continues to rewire the search tree and updates the solution (line 1, Algorithm 1). If a solution regarding current $\mathcal{G}$ is not found, the inner loop reports failure and terminates. In case of early termination of the inner loop with remaining time, the search graph $\mathcal{G}$ is extended in the outer loop to enhance potential solution quality (line 1, Algorithm 1).
@@ III-D Search Graph Densification
The ExtendSearchGraph procedure (Algorithm 6) extends the underlying search graph $\mathcal{G}$ to improve both completeness and optimality of the solution. $\mathcal{G}$ is grew to incrementally approximate the problem domain by sampling a random point $v_{rand}$ from $\mathcal{X}$ (line 6, Algorithm 6) and extending some parts of the graph towards $v_{rand}$. First, the closest node in the graph $v_{nearest}$ is found, then a new sample $v_{new}$ is generated by steering $v_{nearest}$ towards a randomly sampled point $v_{rand}$ by distance of $\delta$ (line 6, Algorithm 6). Then, the lazily-evaluated edges among $v_{new}$ and its neighbor nodes are added to $\mathcal{G}$ (line 6-6, Algorithm 6). In this paper, the neighbor nodes are found by searching for nodes that are within radius $r$ of $v_{new}$, which is given in [6] by
@@
$r=\min\left\{2\left(1+\frac{1}{d}\right)^{\frac{1}{d}}\left(\frac{\mu\left(\mathcal{X}_{\text{free }}\right)}{\zeta_{d}}\right)^{\frac{1}{d}}\left(\frac{\log(N)}{N}\right)^{\frac{1}{d}},\delta\right\}$ (4)
@@
where N=$|V|$, $d$ is the dimension of the problem, $\mu\left(\mathcal{X}_{free}\right)$ is the Lebesgue measure of the free configuration space $\mathcal{X}_{free}$, $\zeta_{d}$ is the volume of a unit ball in $\mathbb{R}^{d}$. The lmc- and g-value of each new node are initialized as $\infty$. If $v_{new}$ is able to yield a solution path via one of its neighbor node, its lmc-value will be updated, and the node will be inserted into $\mathcal{Q}$ (line 6, Algorithm 6). Later, the ComputeShortestPath procedure will propagate the new information due to the extension across the graph and search for a better solution for promising nodes.
IV Simulation Results
In this section, we present simulation results comparing LLPT against RRT[6], Informed RRT[8], BIT[19], RRTX[15], and RT-RRT[14] to verify the performance of LLPT. We compare them for static and dynamic path planning problems in $\mathbb{R}^{2}$, $\mathbb{R}^{3}$. The length for each dimension is 30.0. The $\epsilon$ parameter of RRTX is set as 0 for all simulations. BIT samples 150 nodes per batch regardless of problem dimension. The Euclidean norm is used as the weight and the heuristic functions for all problems. The edge resolution is set as 0.02 for collision check. The saturate parameter $\delta$, equivalent to the maximum edge length of the search graph, is set as 2 for $\mathbb{R}^{2}$ and 3 for $\mathbb{R}^{3}$. The simulations were developed in Python 3.0 and ran on a Macbook with an M2 chip and 16GB RAM.
Figure 1: The x-y plane of the obstacle configurations (represented by black regions) of the static planning problems. The red dot and the blue dot denotes the start and the goal configuration, respectively.
IV-A Static Planning Problems
The planners were tested on two abstract problems with different obstacle configurations. The x-y plane of the obstacle configurations is visualized in Fig.1. The start and goal are located at $[10,15,15,...,15]^{T}$ and $[20,15,15,...,15]^{T}$ for the problem illustrated in Fig.1(a), and $[8,8,8,...,8]^{T}$ and $[23,23,23,...,23]^{T}$ for the problem illustrated in Fig.1(b). The evaluation parameter $\alpha$ of LLPT is set as 1. Each planner was tested with 50 different pseudo-random seeds.
(a) $\mathbb{R}^{2}$, Fig.1(a)
(b) $\mathbb{R}^{2}$, Fig.1(b)
(c) $\mathbb{R}^{3}$, Fig.1(a)
(d) $\mathbb{R}^{3}$, Fig.1(b)

### v3

@@ 5       $V\xleftarrow{+}\{v_{new}\}$;
8       if free-space local trajectory exists between $v$ and $u$ then
9             $E\xleftarrow{+}\{e(v_{new},u),e(u,v_{new})\}$;
@@ 10             UpdateNode($v,u$);
12
13
@@ III-B Informed Rewiring Cascade
LLPT dynamically adjusts the solution path by rewiring $\mathcal{T}$ when the edge weights of $\mathcal{G}$ change. Drawing inspiration from RRTX [36] and RRT [7], LLPT employs two cost-to-goal estimates, g($v$) and lmc($v$), for each node to detect local inconsistencies arising from graph modifications. A node $v$ is called consistent if g($v$) = lmc($v$)<absent<$\infty$< ∞, and inconsistent otherwise. An inconsistent node is further categorized as locally overconsistent if lmc($v$) $<$ g($v$), or locally underconsistent if lmc($v$) $>$ g($v$) or lmc($v$) = g($v$) $\infty$. The g-value is the accumulated cost-to-goal by traversing the previous search tree, while the lmc-value is the minimal cost-to-goal based on the up-to-date graph topology and is potentially better informed than the g-value. The lmc-value satisfies the following relationship when all nodes are locally consistent:
@@
Modifications on edge weights potentially trigger changes in the lmc-value of some nodes and render these nodes inconsistent. To minimally rewire the inconsistent nodes while guaranteeing solution completeness and optimality, we design the rewiring cascade based on the following observations.
Observation 1.
If the cost of the solution path from a node to the goal increases, the cost-to-goal of all its descendants (if it has any) in $\mathcal{T}$ will implicitly increase by the exact change. The node and its descendants may find a shorter path to the goal among their neighboring nodes belonging to $\mathcal{T}$. Therefore, the node and its descendants should be rewired to guarantee solution completeness after rewiring $\mathcal{T}$.
Observation 2.
If the cost of the solution path from a node to the goal decreases, the cost-to-goal of all its descendants (if it has any) in $\mathcal{T}$ will implicitly decrease by the exact change. The node and its descendants may provide a shortcut to their neighboring nodes. Therefore, all the neighboring nodes should be rewired to guarantee solution optimality.
LLPT collects and stores the inconsistent nodes in a priority queue $\mathcal{Q}$. The priority of each node in the queue is based on the order of its associated keys. If a node $v$ becomes underconsistent, which happens when its solution path to the goal is evaluated to be in collision, the node and all its leaf nodes should be rewired to guarantee solution completeness according to Observation 1. LLPT will use the PropagateCostToLeave procedure (Algorithm 3) to propagate the cost ascendance to all its leaf nodes. Afterward, both $v$ and its leaf nodes will be added to the priority queue $\mathcal{Q}$. On the other hand, if a node $v$ becomes overconsistent, which can happen when some of its outgoing edges have decreasing weight or when $\mathcal{G}$ is extended, $v$ will be added to $\mathcal{Q}$ to trigger rewiring among its neighboring nodes to ensure solution optimality according to Observation 2.
The ComputeShortestPath procedure, which is provided in Algorithm 4, is called to make the spanning tree consistent again by operating on the inconsistent nodes stored in $\mathcal{Q}$ iteratively. For an underconsistent node, ComputeShortestPath iteratively searches among its neighbor nodes, and the underconsistent node is rewired to a new parent node if the new parent node results in a better cost-to-goal (line 4-4, Algorithm 4). For an overconsistent node, its local inconsistency is disseminated to its neighbors. Specifically, if $v$ emerges as a better parent node for a neighbor node, the neighbor’s parent is updated to $v$, and the neighbor will be included in $\mathcal{Q}$ to propagate the cost-to-goal reduction (line 4-4, Algorithm 4).
@@
where $\dot{<}$ is a lexicographical comparator with $k(v)\dot{<}k(u)$ if either $k_{1}(v)<k_{2}(u)$ or $k_{1}(v)=k_{1}(u)\land k_{1}(v)<k_{2}(u)$. The node in $\mathcal{Q}$ with the smallest key is regarded as the most promising one, which is popped out from $\mathcal{Q}$ and processed first until there is no node with smaller key than $v_{start}$ and $v_{start}$ has found a solution path to $v_{goal}$ (line 1, Algorithm 4). Since $v_{start}$ is updated as the robot moves, $h(v_{start},v)$ is changing and thus the key used by the priority queue. To avoid having to reorder the queue, we adopt the key updating strategy as D Lite[5], where a constant $k_{m}$ (initialized as 0) that is added up with the cost of the robot movement at each replanning iteration is used to mitigate the priority inconsistency between the out-of-date priority in $\mathcal{Q}$ and the priority of a newly-inserted vertex (line 1, Algorithm 1). The key of a newly-inserted vertex $k(v)$ is calculated by $k_{1}(v)=\min(lmc(v),g(v))+h(v_{start},v)+k_{m}$ and $k_{2}(v)=\min(lmc(v),g(v))$. For each node popped from the top of $\mathcal{Q}$, its key will be recalculated, and the node will be reinserted into $\mathcal{Q}$ with updated priority if its old key is smaller than the new one (line 4-4, Algorithm 4). This assures the efficiency of the queue order.
@@ III-C Algorithmic framework
The proposed algorithm is provided in Algorithm 1 with its major sub-routines outlined in Algorithms 2-6. Notations such as $A\xleftarrow{+}B$ represents $A\leftarrow A\cup B$, and $A\xleftarrow{-}B$ represents $A\leftarrow A\backslash B$. Set cardinality is denoted by ||$\cdot$|| ⋅ |. $\textnormal{{Child}}(v)$ and $\textnormal{{Parent}}(v)$ retrieve child nodes and the parent of node $v$ in $\mathcal{T}$, respectively. $N(v)$ returns the neighboring nodes of $v$ in $\mathcal{G}$. The priority queue $\mathcal{Q}$ stores nodes sorted in ascending order of their keys. $\mathcal{Q}.\textnormal{{update}}(v)$ inserts or re-orders node $v$ based on its key value. $\mathcal{Q}.\textnormal{{pop}}()$ pops and returns the node with the highest priority in $\mathcal{Q}$. $\mathcal{Q}.\textnormal{{topkey}}$ returns the key of the vertex with the highest priority in $\mathcal{Q}$. The function $\textnormal{{MakeParentOf}}(v,u)$ establishes a child-parent relationship between $v,u$ and adds $v$ to $\mathcal{T}$. $\textnormal{{RemoveFromTree}}(v)$ eliminates the child-parent relationship between $v$ and its parent node and removes $v$ from $\mathcal{T}$. The comparison operator with $\infty$ is defined such that $x<\infty$ is true for any $x\neq\infty$ and false otherwise.
@@
LLPT initiates by setting the goal node $v_{goal}$ with lmc($v_{goal}$)=g($v_{goal}$)=0 and the start node $v_{start}$ with lmc($v_{start}$)=g($v_{start}$)$\infty$. The replanning procedure (line 1-1, Algorithm 1) iterates indefinitely at a user-defined frequency until the robot reaches the goal configuration. At the start of each replanning cycle, the search graph $\mathcal{G}$ is refreshed with lazily-computed edges (line 1, Algorithm 1). For edges with updated weights, their starting nodes are rewired using UpdateNode (line 1-1, Algorithm 1). Each replanning cycle comprises two primary loops: the outer loop (line 1-1, Algorithm 1) and the inner loop (line 1-1, Algorithm 1). The inner loop serves as the main search process, rewiring the search tree $\mathcal{T}$ and identifying the shortest collision-free path. This loop alternates between rewiring $\mathcal{T}$ to compute the shortest path and evaluating the edges along the shortest path until a collision-free shortest path is discovered. The procedure EvaluateEdge evaluates the first $\alpha$ unevaluated edges of the shortest path, where $\alpha\in\mathbb{N}^{*}$ is a user-specified parameter to balance the competing computational costs of edge evaluation and rewiring, and returns the collided nodes. For the collided nodes, they are removed from $\mathcal{T}$. Their lmc-value is set to infinity, and they are then inserted into $\mathcal{Q}$ (line 3-3, Algorithm 3). The cost-to-goal increase information is iteratively propagated to their descendants (line 3-3, Algorithm 3). The inner loop continues to rewire the search tree and updates the solution (line 1, Algorithm 1). If no solution for the current $\mathcal{G}$ is found, the inner loop reports failure and terminates (line 1, Algorithm 1). When the inner loop terminates with remaining planning time, the search graph $\mathcal{G}$ is densified in the outer loop to enhance potential solution quality (line 1, Algorithm 1).
@@ III-D Search Graph Densification
The ExtendSearchGraph procedure (Algorithm 6) extends the underlying search graph $\mathcal{G}$ to improve both completeness and optimality of the solution. $\mathcal{G}$ is growing to incrementally approximate the problem domain by sampling a random point $v_{rand}$ from $\mathcal{X}$ (line 6, Algorithm 6) and extending some parts of the graph towards $v_{rand}$. First, the closest node in the graph $v_{nearest}$ is found (line 6, Algorithm 6). Then, $v_{rand}$ is brought $\delta$-closer to $v_{nearest}$, where the values of $\delta$ is problem-specific. The lazily-evaluated edges among $v_{new}$ and its neighbor nodes are added to $\mathcal{G}$ (line 6-6, Algorithm 6). In this paper, the neighbor nodes are found by searching for nodes that are within radius $r$ of $v_{new}$, which is given in [6] by
@@
$r=\gamma_{s}2\left(1+\frac{1}{d}\right)^{\frac{1}{d}}\left(\frac{\mu\left(\mathcal{X}_{\text{free }}\right)}{\zeta_{d}}\right)^{\frac{1}{d}}\left(\frac{\log(N)}{N}\right)^{\frac{1}{d}}$ (4)
@@
where $\gamma_{s}$ is the tuning parameter, N=$|V|$, $d$ is the dimension of the problem, $\mu\left(\mathcal{X}_{free}\right)$ is the Lebesgue measure of the free configuration space $\mathcal{X}_{free}$, $\zeta_{d}$ is the volume of a unit ball in $\mathbb{R}^{d}$. The lmc- and g-value of each new node are initialized as $\infty$. If $v_{new}$ is able to yield a solution path via one of its neighbor nodes, its lmc-value will be updated, and the node will be inserted into $\mathcal{Q}$ (line 6, Algorithm 6). Later, the ComputeShortestPath procedure will propagate the new information due to the extension across the graph and search for a better solution for promising nodes.
IV Analysis
In this section, we provide theoretical properties concerning the probabilistic completeness and asymptotic optimality of LLPT. First, the following theorem is given:
Theorem 1.
Upon the termination of the ComputeShortestPath procedure, the optimal path from $v_{start}$ to $v_{goal}$ in $\mathcal{G}$ regarding the lazy weight function $\bar{w}$ is returned, i.e., $\pi=\arg\min_{\Pi\in\mathcal{G}}\sum_{e\in\pi}\bar{w}(e)$, where $\Pi$ is the set of all finite cost paths connecting $v_{start}$ to $v_{goal}$ in $\mathcal{G}$. Otherwise, if upon the termination there is no solution path that connects $v_{start}$ and $v_{goal}$, there is no solution regarding the current search graph.
Proof.
See Appendix -A. ∎
Theorem 2.
The LLPT algorithm finds the optimal solution with respect to the current search graph with fully-evaluated edges if there exists one when the inner search loop (line 1-1, Algorithm 1) terminates.
Proof.
The inner search loop terminates when (1) the ComputeShortestPath procedure terminates without a solution path that connects $v_{start}$ and $v_{goal}$, and (2) the ComputeShortestPath procedure terminates with a solution path $\pi$ which is fully evaluated and is found to be collision-free. For the first case, as proven by Theorem 1, there is no solution regarding the lazy weight function $\bar{w}$. As $\bar{w}(e)<w(e)$ for $\forall e\in E$, the number of collision-free edges the fully-evaluated search graph has must be no more than that of the lazy-evaluated search graph. Hence, if the lazy-valuated search graph does not yield a solution, the fully-evaluated one can not yield one either. For the second case, as proven by Theorem 1, $\pi$ is the optimal path from $v_{start}$ to $v_{goal}$ regarding the lazy weight function $\bar{w}$. If $\pi\in E_{eval}$, we have ${w}(\pi)=\sum_{e\in\pi}{w}(e)=\sum_{e\in\pi}\bar{w}(e)\leq\sum_{e\in\pi^{*}}\bar{w}(e)\leq\sum_{e\in\pi^{*}}{w}(e)$, where $\pi^{*}$ is the optimal solution path w.r.t. the current search graph with fully-evaluated edges. Therefore, $\pi$ is the optimal path w.r.t. the current search graph. ∎
Theorem 3 (Asymptotical Optimality).
The LLPT algorithm is asymptotically optimal as planning time $t\rightarrow\infty$, that is, LLPT will return the optimal solution path w.r.t. the known environment and the cost function $w(\cdot)$ given infinite planning time.
Proof.
The LLPT algorithm will extend the underlying search graph infinitely by the ExtendSearchGraph procedure when there is planning time remaining. The ExtendSearchGraph procedure will create a search graph $\mathcal{G}$ of which the RRG whose edge weight is evaluated by the lazy weight function $\bar{w}$ is a subgraph, which is proven to be surely containing an optimal solution with finite cost for any non-negative cost function as long as there exists one when the number of nodes is approaching $\infty$ [6]. As proven by Theorem 2, the ComputeShorestPath procedure will return the optimal solution w.r.t. $\mathcal{G}$ with fully-evaluated edges, or equivalently, the RRG with fully-evaluated edges. Therefore, LLPT will eventually converge to the optimal solution as long as there exists one. ∎
V Simulation and Experiment Results
(a) $\mathbb{SE}(3)$ Cubicles
(b) $\mathbb{SE}(3)$ Piano Mover
(c) $\mathbb{R}^{7}$ Emika Franka Robot Arm
Figure 3: Depictions of rigid-body motion planning problems from the OMPL.app and PyBullet.
(a) Cubicles, $\mathbb{SE}(3)$
(b) Piano Mover, $\mathbb{SE}(3)$
(c) Manipulation, $\mathbb{R}^{7}$